OpenStudy (anonymous):

Heres another one :D An alloy of tin and copper contains 16% tin. How much tin (to the nearest tenth) must be added to make 820 pounds of alloy that is 18% tin?

OpenStudy (dumbcow):

Alloy Tin Before x .16x After 820 820*.18 =147.6 Since we are adding only tin, the difference in weight should be the same 820-x = 147.6 - .16x .84x = 672.4 x = 800.48 difference = 820 - 800.48 = 19.52

OpenStudy (ar43r):

nice solution!